# Conceptualization, Formal analysis, Investigation, Methodology, Software, Data curation, Validation, Visualization, Writing –original draft, Writing – review & editing

### v3

@@ Abstract
Accurate multi-sensor calibration is essential for deploying robust perception systems in applications such as autonomous driving, robotics, and intelligent transportation. Existing LiDAR-camera calibration methods often rely on manually placed targets, preliminary parameter estimates, or intensive data preprocessing, limiting their scalability and adaptability in real-world settings. In this work, we propose a fully automatic, targetless, and online calibration framework, CalibRefine, which directly processes raw LiDAR point clouds and camera images. Our approach is divided into four stages: (1) a Common Feature Discriminator that trains on automatically detected objects–using relative positions, appearance embeddings, and semantic classes–to generate reliable LiDAR-camera correspondences, (2) a coarse homography-based calibration, (3) an iterative refinement to incrementally improve alignment as additional data frames become available, and (4) an attention-based refinement that addresses non-planar distortions by leveraging a Vision Transformer and cross-attention mechanisms. Through extensive experiments on two urban traffic datasets, we show that CalibRefine delivers high-precision calibration results with minimal human involvement, outperforming state-of-the-art targetless methods and remaining competitive with, or surpassing, manually tuned baselines. Our findings highlight how robust object-level feature matching, together with iterative and self-supervised attention-based adjustments, enables consistent sensor fusion in complex, real-world conditions without requiring ground-truth calibration matrices or elaborate data preprocessing.
keywords:
[…]
1 Introduction
Reliable and accurate environment perception is becoming increasingly pivotal for advanced applications such as autonomous driving, robotics, and intelligent transportation systems. High-quality perception is essential for making well-informed decisions and ensuring safe, efficient operations. However, relying on a single sensor often presents significant challenges (Lv et al., 2021; Sengupta et al., 2022). For instance, cameras provide rich color and texture information (Berrio et al., 2021) but are highly sensitive to lighting conditions and can struggle with depth estimation, especially in poorly illuminated or rapidly changing environments (An et al., 2024). Conversely, LiDAR sensors offer precise 3D geometric information (Zhang and Rajan, 2022) and are robust to illumination variations, yet they can be more expensive and may suffer from reduced performance under adverse weather conditions (Northrop et al., 2022). Consequently, multi-sensor fusion has emerged as a practical strategy to leverage the complementary strengths of various sensing modalities (Cui et al., 2021; Ye et al., 2021; Cheng et al., 2024). Among these, camera–LiDAR fusion stands out for its ability to integrate high-resolution visual details with accurate depth measurements (Jiao et al., 2023). Despite the evident benefits, the success of sensor fusion fundamentally hinges on accurate inter-sensor calibration (Duan et al., 2023; Yoon et al., 2021). In fact, any imprecision or unreliability in this calibration process can severely degrade the performance of downstream perception tasks, leading to erroneous object detection, flawed localization, and ultimately compromised system reliability.
Calibration, as a critical requirement in ensuring accurate data fusion from multiple sensors, is commonly divided into three categories: intrinsic, extrinsic, and temporal calibration (Peršić et al., 2021). Intrinsic calibration focuses on determining each sensor’s internal parameters and is often provided by the manufacturer or derived from well-established procedures (Lv et al., 2022; Domhof et al., 2021). Temporal calibration estimates the constant offset between measurement instants and timestamps (Yuan et al., 2022), typically caused by jitter (random message delays), skew (differences in clock rates), sampling-frequency mismatches, or communication delays (Rehder et al., 2016). In practice, temporal calibration can be addressed either through hardware synchronization—where all sensors share a common clock—or through software-based synchronization approaches, such as approximate time synchronization (Yeong et al., 2021), which can align data by analyzing timestamp differences. Although intrinsic and temporal calibrations are vital, they are generally managed separately under controlled conditions and can often achieve satisfactory results. Consequently, in multi-sensor systems, the primary challenge shifts to extrinsic calibration—also referred to as spatial calibration or sensor registration—which involves determining the relative spatial transformations among sensor coordinates (Domhof et al., 2019; Qiu et al., 2023; Cheng et al., 2023). In the case of LiDAR–camera systems, extrinsic calibration fundamentally involves collecting corresponding points in both the LiDAR and camera coordinate frames (Cheng et al., 2023; Li et al., 2023) and using these correspondences to compute the rigid-body transformation matrix that maps one coordinate system to the other (Pandey et al., 2012; Park et al., 2020).
Depending on how these point correspondences are obtained, extrinsic calibration methods can be classified along several axes. For instance, target-based methods rely on carefully designed calibration artifacts such as checkerboards or fiducial markers to ensure accurate feature correspondence (Xiao et al., 2024). While these can deliver high-precision correspondences, they demand considerable effort in terms of setup, controlled environment preparation (Cheng and Cao, 2023), and the repetitive placement of targets (Cheng et al., 2023)—factors that are often impractical or costly in real-world applications. By contrast, targetless approaches dispense with dedicated calibration objects and instead exploit naturally occurring features in the environment (Cheng and Cao, 2023; Chen et al., 2022). This eliminates the overhead of managing physical targets and can be more adaptable to diverse scenarios, though it may pose challenges if the environment offers insufficient or indistinct features. A second dimension pertains to whether correspondence collection is manual or automatic. Manual calibration typically requires an operator to identify or verify matching points across sensors, potentially achieving high accuracy at the expense of significant time and labor (Yin et al., 2023). This approach lacks scalability, especially in large-scale or real-time deployments. Automatic calibration, on the other hand, autonomously extracts and matches features (Sun et al., 2022b; Zhang et al., 2023a), thus minimizing human involvement (Li et al., 2023). Although it hinges on the robustness of detection and matching algorithms, its reduced labor demands and potential for continuous operation make it highly appealing for applications where speed and scalability are paramount. Finally, calibration can be performed offline or online (Schneider et al., 2017; Zhu et al., 2023; Shi et al., 2020). Offline processes usually operate on batched data, allowing for extensive optimization and generally delivering precise results. However, offline methods cannot adapt to real-time sensor shifts, environmental changes, or hardware reconfigurations. In contrast, online calibration continuously updates transformation parameters as new sensor data arrives, accommodating dynamic conditions but introducing additional computational and algorithmic complexity.
Given these trade-offs, a fully automatic, targetless, and online calibration paradigm combines the most desirable attributes—removing cumbersome calibration objects, eliminating the need for human intervention, and adapting in real time to changing environments. Yet achieving all three simultaneously presents significant challenges, particularly in feature detection and correspondence matching. Nevertheless, recent advances in deep learning—especially in robust feature extraction and cross-modal matching—have opened up new pathways toward achieving this.
Current automatic, targetless, and online calibration methods for LiDAR–camera systems remain scarce, and those that do exist often exhibit significant drawbacks. Motion-based methods (Petek et al., 2024; Park et al., 2020; Sun et al., 2022a), for instance, either require extra hardware—such as camera- and LiDAR-based odometry systems—or impose constraints ill-suited to many real-world applications. Hand–eye calibration (Ou et al., 2023; Sun et al., 2022a), although well-established, is notoriously cumbersome and time-consuming, demanding multiple sensor poses that are difficult to obtain accurately in real-world dynamic scenarios where objects are in motion. Additionally, hand–eye calibration is ill-suited for static sensor setups, such as those mounted on vehicles or fixed traffic infrastructure. Edge-based calibration approaches (Li et al., 2024; Yuan et al., 2021; Yin et al., 2023; Zhang and Rajan, 2022; Zhang et al., 2023b) attempt to align object edges detected by the sensors, but matching edges is inherently difficult because object boundaries are often arbitrary and appear markedly different between LiDAR and camera modalities. Mutual information–based methods (Pandey et al., 2012; Taylor and Nieto, 2013; Koide et al., 2023) that often rely on reflectance intensities are similarly unreliable because LiDAR reflectance varies across materials and surfaces, while camera pixel intensities are heavily influenced by lighting conditions. Deep learning methods that directly regress calibration parameters—exemplified by RegNet (Schneider et al., 2017) and its variants (Zhu et al., 2023; Lv et al., 2021; Xiao et al., 2024; Duan et al., 2024; Shi et al., 2020; Iyer et al., 2018; Shang and Hu, 2022)—need an initial calibration to project LiDAR data into the image plane, resulting in a "projected LiDAR depth map". While innovative, this approach has several limitations. First, it relies on an initial calibration, which is typically obtained through manual intervention or empirical estimates based on the sensor setup, reducing generalizability. Second, the use of projected LiDAR depth maps heavily emphasizes image data, as LiDAR point clouds are sparse compared to image pixels. This results in extracted features being dominated by image characteristics, effectively sidelining valuable LiDAR information. Lastly, directly regressing calibration matrices is akin to solving an unconstrained task within constrained optimization, which is computationally challenging for current neural networks. Such models also suffer from limited generalization and are incompatible with real-time requirements due to their computational overhead. Moreover, most existing methods fail to fully leverage the already advances in object detection and feature extraction achieved in the respective domains of LiDAR and camera data processing. While some approaches utilize semantic segmentation (Sun et al., 2022a; Luo et al., 2024; Jiang et al., 2021), they often focus solely on image-based segmentation, which is computationally expensive and offers no significant advantage over object detection for calibration tasks.
@@ 1 Introduction
In light of these limitations, we present a fully automatic, targetless, online calibration framework (Fig. 1), CalibRefine, that utilizes raw, unmodified LiDAR point clouds and camera images, avoiding any prior knowledge or preprocessing, such as initial calibration or projected depth maps. Instead of relying on ground-truth calibration matrices, our method employs a supervised phase for coarse calibration (using only object-level labels) followed by a self-supervised post-refinement stage. First, we leverage robust object detection algorithms—YOLOv8 (Jocher, 2023) for camera images and an octree change detection (Strawlab, 2023)-plus-DBSCAN pipeline for LiDAR data—to detect and extract individual objects from each sensor’s output. A Common Feature Discriminator is then trained to match these object instances by learning more comprehensive and distinctive object-level features, going beyond simple edge features, as shown in Fig. 2. It leverages three complementary features: (1) relative positions within each sensor’s frame, (2) appearance embeddings, extracted using ResNet (He et al., 2015) for camera data and PointNet++ (Qi et al., 2017) for LiDAR data, and (3) object classification information, to offer robust cross-sensor object matching capabilities. These matched objects provide a pool of reliable point correspondences, from which we compute a coarse calibration matrix using a homography-based approach. Recognizing that the Common Feature Discriminator may not achieve perfect accuracy in object matching, we enhance calibration precision through two online refinement processes. The iterative optimization-based refinement projects LiDAR points onto the camera plane, matches them with camera points using a greedy bipartite graph algorithm, and refines the calibration matrix every N frames based on reprojection error improvements. The attention-based refinement utilizes a Vision Transformer (Dosovitskiy et al., 2021) to perceive global distortion features from images, effectively mitigating the limitations of homography calibration caused by non-planar surfaces and depth variations. Furthermore, it incorporates a cross-attention network to compute weighted interactions between image pixels (queries) and LiDAR points (keys and values), establishing robust spatial correspondences and achieving further refinement of the calibration matrix. Crucially, our framework bypasses the pitfalls of direct matrix regression and the need for projected LiDAR maps, and eliminates the reliance on heuristic preprocessing or manually labeled calibration matrices, offering a more principled, data-driven pipeline that is both computationally efficient and adaptable in real-time. By integrating domain-specific mature object detection methods, a reliable discriminator to identify cross-sensor correspondences, and dual-stage refinement, our approach bridges the existing research gap, achieving a stable and accurate LiDAR–camera calibration that is truly automatic, targetless, and online. Our contributions are summarized as follows:
@@
    Coarse-to-Fine Calibration Strategy with Dual Refinement Processes: The framework adopts a two-stage calibration approach, combining a homography-based coarse calibration with iterative refinement and attention-based refinement methods. These processes improve calibration accuracy in real-time, addressing challenges such as non-planar surfaces and dynamic conditions.

### v4

# CalibRefine: Deep Learning-Based Online Automatic Targetless LiDAR–Camera Calibration with Iterative and Attention-Driven Post-Refinement

@@ Abstract
Accurate multi-sensor calibration is essential for deploying robust perception systems in applications such as autonomous driving and intelligent transportation. Existing LiDAR-camera calibration methods often rely on manually placed targets, preliminary parameter estimates, or intensive data preprocessing, limiting their scalability and adaptability in real-world settings. In this work, we propose a fully automatic, targetless, and online calibration framework, CalibRefine, which directly processes raw LiDAR point clouds and camera images. Our approach is divided into four stages: (1) a Common Feature Discriminator that trains on automatically detected objects–using relative positions, appearance embeddings, and semantic classes–to generate reliable LiDAR-camera correspondences, (2) a coarse homography-based calibration, (3) an iterative refinement to incrementally improve alignment as additional data frames become available, and (4) an attention-based refinement that addresses non-planar distortions by leveraging a Vision Transformer and cross-attention mechanisms. Through extensive experiments on two urban traffic datasets, we show that CalibRefine delivers high-precision calibration results with minimal human involvement, outperforming state-of-the-art targetless methods and remaining competitive with, or surpassing, manually tuned baselines. Our findings highlight how robust object-level feature matching, together with iterative and self-supervised attention-based adjustments, enables consistent sensor fusion in complex, real-world conditions without requiring ground-truth calibration matrices or elaborate data preprocessing. Code is available at https://github.com/radar-lab/Lidar_Camera_Automatic_Calibration
Index Terms:
LiDAR-camera calibration, extrinsic calibration, online automatic calibration, sensor fusion, traffic monitoring
I Introduction
Reliable and accurate environment perception is crucial for applications such as autonomous driving, robotics, and intelligent transportation systems, enabling informed decisions and ensuring safe, efficient operations. However, single-sensor perception often encounters inherent limitations [1, 2]. Cameras provide rich visual detail but are sensitive to lighting variations and struggle with depth estimation, especially under poor illumination or dynamic conditions [3, 4]. LiDAR sensors, conversely, provide precise 3D geometric measurements robust to lighting changes, yet can be costly and suffer performance degradation under adverse weather [5, 6]. Consequently, multi-sensor fusion, particularly LiDAR-camera fusion, has gained prominence, integrating visual textures with accurate spatial data [7, 8]. Nonetheless, successful fusion fundamentally relies on accurate sensor calibration, as imprecise calibration severely compromises downstream perception accuracy [9, 10].
Calibration procedures generally include intrinsic, extrinsic, and temporal calibration [11]. While intrinsic calibration (determining internal sensor parameters) and temporal calibration (synchronizing timestamps) typically follow standardized practices and achieve reliable results [12, 13, 14], extrinsic calibration—also known as spatial calibration—remains challenging. Extrinsic calibration seeks to identify spatial transformations between sensor coordinate systems, usually by establishing correspondences between matched points [15, 16]. Extrinsic calibration methods vary primarily by how these correspondences are established: target-based approaches utilize dedicated calibration artifacts (e.g., checkerboards or markers) for precise correspondences, yet they require elaborate setup and are impractical for real-world scenarios [17, 18]; targetless methods instead use natural scene features, eliminating cumbersome preparations but posing challenges when suitable features are scarce or indistinct [18, 19]. Methods also differ regarding human intervention (manual versus automatic) and operational mode (offline versus online). Manual methods, although precise, involve significant human effort and thus lack scalability [20]. Automatic methods autonomously establish feature correspondences, minimizing manual intervention, making them attractive for scalable and continuous operation [21, 22]. Offline calibration methods rely on batch data and extensive optimizations but fail to adapt to real-time sensor shifts, environmental changes, or hardware reconfigurations. In contrast, online calibration continuously updates calibration parameters as new data arrives, accommodating dynamic conditions, albeit with increased computational complexity [23, 24, 25].
Given these trade-offs, a fully automatic, targetless, and online calibration paradigm combines the most desirable attributes—removing cumbersome calibration objects, eliminating the need for human intervention, and adapting in real time to changing environments. However, despite its appeal, such a paradigm remains highly challenging, and existing approaches in this category exhibit various limitations. Motion-based methods often rely on additional hardware or specific sensor motion constraints, limiting real-world applicability [26, 27]. Hand-eye calibration, while classical, demands multiple accurate sensor poses, rendering it unsuitable for static or dynamically constrained setups [28]. Edge-based methods suffers from unreliable edge matching across modalities, since object boundaries differ significantly across LiDAR and camera data [29, 30, 5]. Mutual information-based methods are similarly unreliable due to varying LiDAR reflectance and camera illumination sensitivities [31, 32]. Recent deep learning–based methods directly regress calibration parameters (e.g., RegNet [23] and its variants [24, 1, 17, 33, 25, 34, 35]) but require initial manual calibrationsor projected depth maps, and suffer from limited generalization and computational overhead, making them less suitable for real-time applications [25, 17].
@@ I Introduction
To address these challenges, we propose CalibRefine, a fully automatic, targetless, online LiDAR–camera calibration framework. Our method directly processes raw LiDAR point clouds and camera images without initial calibration matrices or complex preprocessing. First, we leverage robust object detection algorithms—YOLOv8 for camera data and an octree-based DBSCAN approach for LiDAR—to identify individual objects. A novel Common Feature Discriminator then matches these cross-sensor object instances by learning relative positions, appearance embeddings (using ResNet [36] for camera images and PointNet++ [37] for LiDAR data), and semantic class information, forming reliable cross-sensor correspondences. Recognizing potential inaccuracies from initial matching, we further enhance calibration through two online refinement stages. The iterative refinement incrementally optimizes calibration using accumulated correspondence data, while the attention-based refinement employs a Vision Transformer [38] and cross-attention mechanisms to correct for non-planar distortions and depth variations, further improving calibration accuracy. Crucially, our framework bypasses the pitfalls of direct matrix regression and the need for projected LiDAR maps, and eliminates the reliance on heuristic preprocessing or manually labeled calibration matrices, offering a more principled, data-driven pipeline that is both computationally efficient and adaptable in real-time. By integrating domain-specific mature object detection methods, a reliable discriminator to identify cross-sensor correspondences, and dual-stage refinement, our approach bridges the existing research gap, achieving a stable and accurate LiDAR–camera calibration that is truly automatic, targetless, and online. Our contributions are summarized as follows:
@@
    Coarse-to-Fine Calibration Strategy with Dual Refinement Processes: The framework adopts a two-stage calibration approach, combining a homography-based coarse calibration with iterative refinement and attention-based refinement methods. These processes improve calibration accuracy in real-time, addressing challenges such as dynamic changes, densely distributed correspondences, and non-planar surfaces.